CompDesk, Inc., makes a single model of an ergonomic desk (with chair) for computer usage. The desk is manufactured in building 1, and the chair is manufactured in building 2. Monthly capacities and production levels are as follows:

Building 1 (Desks) Building 2 (Chairs)
Monthly capacity 400 500
Monthly production 400 400

The company will sell a desk only with a chair and can sell 500 desks per month. The units (desk with chair) sell for $279 each and have a variable cost of $118 each.

Required:

a. Is there a bottleneck at CompDesk in Building 1 or Building 2?

Building 1
Building 2

b. CompDesk’s production supervisors state they could increase building 1’s capacity by 100 desks per month by producing desks on the weekend. Producing on the weekend would not affect the sales price. Variable cost per unit would increase by $18 for those produced on the weekend because of the premium paid to labor. Fixed costs would also increase by $18,600 per month.

b-1. Calculate the differential operating profit (loss). (Losses and amounts to be deducted should be indicated with a minus sign.)

b-2. Should CompDesk produce desks on the weekend?

Yes
No

c. Independent of the situation in requirement (b), CompDesk could add additional equipment and workers to building 1, which would increase its capacity by 100 desks per month. This would not affect the sales price or variable cost per unit but would increase fixed costs by $14,300 per month.

c-1. Calculate the differential operating profit (loss). (Losses and amounts to be deducted should be indicated with a minus sign.)

c-2. Should CompDesk add the additional equipment and workers to building 1?

Yes
No
Answer #1

a.Bottleneck is Building 1 since capacity is less than demand

b-1 Differential profit = Additional sales – additional costs

= (279-118-18)*100-18600

= -$4,300

i.e. loss

b-2 No, since loss

c-1 Differential operating profit = (279-118)*100-14300

= $1,800

C-2, Yes, since profit